OpenStudy (anonymous):

Find the axis of the graph of the function. f(x) = 4x^2 + 8x + 7

OpenStudy (amoodarya):

turn it to standard form y=a(x-b)^2 +c so x=b is axis of symmetry

OpenStudy (anonymous):

So would it be -2?

OpenStudy (amoodarya):

4(x^2+2x) +7 =4 ((x+1)^2- 1)+7 =4(x+1)^ +7-3 so x=-1